Supplemental Operations Report 22Apr2023

[title Supplemental Operations Report – April 22th]

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation.

Power system:
Solar: nominalSOC Last 24 hours: Max 100%; Min 61 %; Avg 78.4%.
VDC Last 24 hours: Max 59.12V; Min 46.35; Avg 51.61V.Generator run time: about 12am – 6am. Not used on April 15th.

Propane Reading Station Tank: 77%
Propane Reading Director Tank: 73%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 77%

Water (Static Tank) – 366 gallons
Water in GreenHab – 117 gallons
Water (Outpost tank) – 442 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: Yes, to Zubrin’s Head and back.
Hours: 190.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged. All corroded battery wires were received and replaced. Sojourner and Spirit need one additional short battery wire to be replaced. Parts are in transit.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. Picked up in Loa after repair, its alternator was replaced. To Hanksville for water runs and supplies pick up.

Crew Car used and why, where? Yes. From Grand Junction by crew. To Hanksville for water runs and supplies.

General notes and comments: Hab Car alternator was replaced.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Nominal

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Nominal. Hab heater was offline several times this week due to strong north winds. Current work is nominal after restart.

Summary of GreenHab Operations: Nominal.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation .

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Shade exterior painting completed. It will require a second coat in a month or so.
Old wood from the clothesline deck is in the process of being burned.
Director’s trailer:
1) Heater went offline after strong winds. The control board displayed a fault code (single red light flashing) – limit switch/airflow problem. Unplugging the control board from power resolved the problem.
2) Porch was leveled.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site